OpenStudy (anonymous):

log2^(x^2+8) =log2^(x) + log2^(6) solve for x

OpenStudy (anonymous):

(x^2+8)log 2 = xlog2 +6l0g 2 (x^2+8)log2= (x+6)log 2. x^2-x+2=0 this dosnt give real value of x